# Edinburgh 2025/29

### v1

@@ 3 Determination of Higher Twist and Jet Power Corrections
The fits presented in this section and the next have been produced using the open-source NNPDF framework [NNPDF:2021uiq] based on the same dataset and theory parameters as used in NNPDF4.0 [NNPDF:2021njg] with MHOUs as in Ref. [NNPDF:2024dpb] and aN3LO corrections as in Ref. [NNPDF:2024nan]. Theory predictions have been generated with the pipeline of Ref. [Barontini:2023vmr], which supersedes the tools used for NNPDF4.0. In particular, interpolation grids are generated using PineAPPL [Carrazza:2020gss] with the DGLAP evolution computed using EKO [Candido:2022tld] and DIS observables using the Yadism package [Candido:2024rkr]. For a detailed discusion of the updated theory pipeline see App A of Ref. [NNPDF:2024djq].
@@ 3.1 Determination of Higher Twist
The DIS measurements included in NNPDF4.0 [NNPDF:2021njg] consist of fixed-target NC structure function data from SLAC [Whitlow:1991uw], BCDMS [Benvenuti:1989rh], and NMC [Arneodo:1996kd, Arneodo:1996qe], fixed-target inclusive and dimuon CC cross-sections from CHORUS [Onengut:2005kv] and NuTeV [Goncharov:2001qe, MasonPhD], and collider NC and CC cross-sections from HERA [Abramowicz:2015mha]. In NNPDF4.0 (and indeed all earlier NNPDF global fits) cuts are applied to all DIS data, $Q^{2}>3.49~{\rm GeV}^{2}$, $W^{2}>12.5~{\rm GeV}^{2}$, in order to avoid the kinematic regions where the HT is expected to be significant.
All calculations of DIS processes in NNPDF4.0 include target mass corrections [Schienbein:2007gr], following the implementation of Refs. [Georgi:1976ve, Ball:2008by], heavy quark mass corrections implemented using FONLL [Forte:2010ta, Ball:2011mu, Barontini:2024xgu], fitted charm [Ball:2015tna, Ball:2015dpa, Ball:2016neh], and nuclear corrections and uncertainties, both for heavy nuclei [Ball:2018twp] and deuteron targets [Ball:2020xqw].
@@
The HT determined in this way is shown in Fig. 3.1 for three different sets of PDFs: NNLO, NNLO with MHOUs, and aN3LO with MHOUs. At NNLO we find that the posterior distribution is significantly more constrained than the prior, justifying the choice of prior. For the proton structure functions, the HT correction $H_{2}^{p}$ is constrained to be very small at low and intermediate $x$, consistent with zero within uncertainties, but increases towards larger values of $x$, reaching a peak of about $4$ GeV2. The deuteron HT correction $H_{2}^{d}$ shows a similar pattern, although with slightly larger uncertainties. Note that, in both cases, the low-$x$ region is well constrained by the data. The HT for CC observables, $H_{CC}$, shows a similar pattern, but is less well constrained, particularly at very small $x$ and very large $x$, where there is no CC data. In particular, at $x=0.9$ there is no constraint at all, with the posterior equal to the prior. This said, we see no evidence here for the expectation [Dasgupta:1996hh], based on renormalons, that HT effects for CC processes are larger than for NC processes, or that they have a different shape.
@@
In Fig. 3.3 we show a comparison of the HT for the $F_{2}$ structure function in proton and deuteron determined in our global NNLO fit, with results obtained in an early PDF fit to just SLAC and BCDMS data by Virchaux and Milsztajn [Virchaux:1991xp]. While this fit was only at NLO, it used a similar multiplicative parametrization to our own, and target mass corrections were included, so the comparison is meaningful. Considering the fact that this fit was performed before the advent of NMC and HERA data, and all the technical improvements made over the intermediate thirty years, the agreement is rather impressive, though it should not come as too much of a surprise since it is the same data (SLAC and BCDMS) that dominates the large-$x$, low-$Q^{2}$ region today. Direct comparisons with more recent determinations which use an additive HT parametrization [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip] are more difficult, since the results necessarily depend to some extent on the choice of scale. However for these too, the results for HT correction to $F_{2}^{p}$ are qualitatively consistent with our own, within uncertainties.
@@
Figure 3.3: Comparison of our NNLO $F_{2}$ higher twist to that determined in protons and deuteron by Virchaux and Milsztajn in a NLO PDF fit to SLAC and BCDMS data [Virchaux:1991xp]. They include TMCs, and use the same multiplicative parametrization of the HT, fitting proton and deuteron data separately, so the results are directly comparable to our own.
@@
When we add MHOUs to our NNLO fit, the HT correction is generally reduced in size. The same happens again when we go to aN3LO, where the MHOUs are significantly reduced. This shows that some of the supposed HT being seen at NNLO is in fact due to MHOUs, principally the N3LO perturbative correction to the leading twist contribution. This shows the difficulty in distinguishing phenomenologically between $Q^{2}$ dependence which is power like, and $Q^{2}$ dependence which is merely a power of a logarithm. This effect was also seen when going from NLO to NNLO in earlier attempts to determine HT [Martin:2003sk]. One is forced to conclude that even the HT determined in the aN3LO fit may still be contaminated by a perturbative contribution. Similarly, $H_{2}^{d}$ may be tangled with the deuteron nuclear correction [Ball:2020xqw], and $H_{CC}$ with the nuclear corrections applied to much of the CC data [Ball:2018twp].
@@ 3.2 Determination of Linear Power Corrections to Jets and Dijets
We now turn to the determination of power corrections for jet observables. The jet cross-sections included in the NNPDF4.0 fit are double differential single-inclusive jet data at $\sqrt{s}=8$ TeV and dijet cross-sections at $\sqrt{s}=7$ TeV from ATLAS [ATLAS:2017kux, ATLAS:2013jmu] and CMS [CMS:2016lna, CMS:2012ftr]. No kinematic cuts were applied to these data in NNPDF4.0. An internal cut on the jet $p_{T}$ was applied to the single-jet measurements of CMS in order to avoid points for which $p_{T}\lesssim 70$ GeV, where fixed-order calculations are not reliable. This cut is not necessary for ATLAS single-jet data since the lowest $p_{T}$ bin starts at $70$ GeV. Finally, no additional cuts are applied to dijet data, as the lowest value of the hard scale is one order of magnitude larger than that for the single-inclusive jet data. These considerations will become relevant in the discussion below.
@@
In Fig. 3.4 we present the posterior distribution for the power corrections to single inclusive jets and dijets, and in Fig. 3.5 the related kinematic distribution of the data points. We observe two distinct patterns for single-inclusive jets and dijets. Starting with the single-inclusive jets, we find that not only is the prior significantly constrained, but also that the posterior distribution is peaked around $10$ GeV. Further, the posterior shows a weak dependence on the rapidity. It should be noted that the previous study in Ref. [Olness:2009qd] estimated a positive power correction of about $7$ GeV for single-inclusive jet cross-sections, independent of rapidity. Our result is consistent with this estimate. For the dijets, we find only weak evidence for a linear PC in the NNLO fit, although uncertainties remain large, showing that the data provide only weak constraints on the PCs. Note that for the dijet kinematics the bins with lowest invariant mass (and thus most likely to be affected by PCs) are at central rapidity, and even here the lowest invariant mass bin is still at a rather higher scale than the lowest $p_{T}$ scale of the single inclusive jet data.
@@
We evaluate the phenomenological consequences of including the HT corrections and PCs discussed in the previous sections at the level of a new global PDF fit, and discuss the impact on relevant LHC observables.
@@ 4.1 Global PDF fits with Power Corrections
The HT and power corrections discussed above were determined by performing global PDF fits to the NNPDF4.0 dataset, but lowering the kinematic cuts to increase sensitivity to the HT. However, we do not suggest performing a fit with these lowered cuts to determine the PDFs, since such a fit would include data which is significantly affected by HT. Instead, we use the HT and the PCs determined above to give corrections to the theory predictions in a fit with the conventional cuts used in NNPDF4.0 [NNPDF:2021njg], i.e. $Q^{2}>3.49$ GeV2 and $W^{2}>12.5$ GeV2.
@@ 4.2 Fit Quality
We start by discussing the fit quality. In Tab. 4.1 we report the values of the experimental $\chi^{2}$ per data point for the global fit with and without HT and PCs, and for fits with perturbative calculations at NNLO, NNLO+MHOU, and aN3LO+MHOU. Together with the global $\chi^{2}$, values are also displayed for each separate process type. In order to disentangle the effect of the shifts in the theory predictions from the effect of the additional uncertainty due to the theory covariance matrix in eq. 4.2, we performed a series of fits where only the shifts are included in the theory predictions. The corresponding $\chi^{2}$ values are also reported in Tab. 4.1. All sources of correlations are fully accounted for, and the $t_{0}$ method [Ball:2009qv] is used for multiplicative uncertainties. We refer to Ref. [NNPDF:2024nan] for further details on the inclusion of the experimental uncertainties in the $\chi^{2}$.
@@ 4.3 PDFs and Parton Luminosities
In Fig. 4.3, the luminosities in the quark-quark, quark-antiquark, gluon-gluon, and gluon-quark channels are shown at $13$ TeV, as ratios of the fits with HT and PCs to the corresponding NNLO, NNLO+MHOU, and aN3LO+MHOU [NNPDF:2024nan] standard fits. As expected from the PDF plots, the effects of HT and PCs at both high and low invariant masses are modest, well within one sigma, and largely disappear once MHOUs are added, and in the aN3LO fit. In fact the only really significant change due to the inclusion of PCs is in the gluon-gluon channel at invariant masses around the Higgs mass. This is particularly noticeable in the standard NNLO fit: here there is a one percent drop in the gluon-gluon luminosity for invariant masses between $100$ and $250$ GeV, just over two sigma. Once MHOUs are included the significance of this effect is reduced, and the aN3LO corrections reduce the impact to well within one sigma. A related effect is seen in the gluon-quark channel, though here it is never more than one sigma.
@@ 4.4 Predictions for LHC observables
We next analyse the impact of power corrections and higher twist on various totally inclusive LHC observables related to DY and Higgs production, specifically vector boson production, associated production of Higgs with a vector boson, then Higgs production in gluon fusion, and in vector boson fusion (VBF). The predictions use hard cross-sections computed up to N3LO using n3loxs [Baglio:2022wzu] for associated production, ggHiggs [Bonvini:2014jma] for gluon fusion, and proVBFH code [Dreyer:2018qbw] for VBF.
@@ 4.5 Effect on the Strong Coupling
We finally consider the effect of HT and PCs on the extraction of the strong coupling at the mass of the $Z$-boson, $\alpha_{s}(m_{Z})$, from the global dataset. We again use the theory covariance methodology, precisely as in Ref. [Ball:2025xgq], but now including also the HT and PC contributions to the theory covariance matrix. The results are shown in Fig. 4.6.
In the NNLO fit, without MHOU, the HT reduces $\alpha_{s}$ a little, while the jet PCs increase it, in each case by around one standard deviation. When MHOU is added, the best fit value of $\alpha_{s}$ drops significantly, and its uncertainty increases substantially, as found in Ref. [Ball:2025xgq]. Again the HT reduces $\alpha_{s}$, and the PCs increase it, by about one standard deviation. However, at aN3LO the pattern changes: while the aN3LO corrections reduce $\alpha_{s}$ a little, the HT correction now results in a substantial increase (more than a standard deviation). We think that this is because at aN3LO the HT is less contaminated by MHO perturbative corrections to the leading twist, and its effect on $\alpha_{s}$ is thus less easily hidden by the MHOU. The effect of the jet PCs is positive as before, but smaller (the jet cross-sections are still evaluated at NNLO+MHOU in the aN3LO fits). The result is a value of $\alpha_{s}(m_{Z})$ at aN3LO which is very consistent with that obtained at NNLO, but with substantially larger uncertainty.
As we noted in Ref. [Ball:2025xgq], all these results are subject to a further uncertainty due to the implementation of positivity constraints on the PDF fit. Since any PDF positivity constraint necessarily introduces a nongaussianity in the distribution of the PDFs, the closure testing of the methodology only works when the positivity constraints are removed. As can be seen in Fig. 4.6, releasing the positivity constraints results in lower values of $\alpha_{s}$ consistently across all the various determinations.

### v2

@@ 3 Determination of Higher Twist and Jet Power Corrections
The fits presented in this section and the next have been produced using the open-source NNPDF framework [33] based on the same dataset and theory parameters as used in NNPDF4.0 [34] with MHOUs as in Ref. [35] and aN3LO corrections as in Ref. [37]. Theory predictions have been generated with the pipeline of Ref. [39], which supersedes the tools used for NNPDF4.0. In particular, interpolation grids are generated using PineAPPL [50] with the DGLAP evolution computed using EKO [48] and DIS observables using the Yadism package [47]. For a detailed discusion of the updated theory pipeline see App A of Ref. [36].
@@ 3.1 Determination of Higher Twist
The DIS measurements included in NNPDF4.0 [34] consist of fixed-target NC structure function data from SLAC [76], BCDMS [42], and NMC [17, 18], fixed-target inclusive and dimuon CC cross-sections from CHORUS [71] and NuTeV [62, 68], and collider NC and CC cross-sections from HERA [9]. In NNPDF4.0 (and indeed all earlier NNPDF global fits) cuts are applied to all DIS data, $Q^{2}>3.49~{\rm GeV}^{2}$, $W^{2}>12.5~{\rm GeV}^{2}$, in order to avoid the kinematic regions where the HT is expected to be significant.
All calculations of DIS processes in NNPDF4.0 include target mass corrections [72], following the implementation of Refs. [61, 27], heavy quark mass corrections implemented using FONLL [60, 23, 40], fitted charm [22, 25, 21], and nuclear corrections and uncertainties, both for heavy nuclei [30] and deuteron targets [31].
@@
The HT determined in this way is shown in Fig. 3.1 for three different sets of PDFs: NNLO, NNLO with MHOUs, and aN3LO with MHOUs. At NNLO we find that the posterior distribution is significantly more constrained than the prior, justifying the choice of prior. For the proton structure functions, the HT correction $H_{2}^{p}$ is constrained to be very small at low and intermediate $x$, consistent with zero within uncertainties, but increases towards larger values of $x$, reaching a peak of about $4$ GeV2. The deuteron HT correction $H_{2}^{d}$ shows a similar pattern, although with slightly larger uncertainties. Note that, in both cases, the low-$x$ region is well constrained by the data. The HT for CC observables, $H_{CC}$, shows a similar pattern, but is less well constrained, particularly at very small $x$ and very large $x$, where there is no CC data. In particular, at $x=0.9$ there is no constraint at all, with the posterior equal to the prior. This said, we see no evidence here for the expectation [55], based on renormalons, that HT effects for CC processes are larger than for NC processes, or that they have a different shape (although it should be remembered that in our analysis sum rule constraints have been imposed only at leading twist).
@@
In Fig. 3.3 we show a comparison of the HT for the $F_{2}$ structure function in proton and deuteron determined in our global NNLO fit, with results obtained in an early PDF fit to just SLAC and BCDMS data by Virchaux and Milsztajn [74]. While this fit was only at NLO, it used a similar multiplicative parametrization to our own, and target mass corrections were included, so the comparison is meaningful. Considering the fact that this fit was performed before the advent of NMC and HERA data, and all the technical improvements made over the intermediate thirty years, the agreement is rather impressive, though it should not come as too much of a surprise since it is the same data (SLAC and BCDMS) that dominates the large-$x$, low-$Q^{2}$ region today. Direct comparisons with more recent determinations which use an additive HT parametrization [15, 13, 43, 11, 10, 14] are more difficult, since the results necessarily depend to some extent on the choice of scale. However for these too, the results for HT correction to $F_{2}^{p}$ are qualitatively consistent with our own, within uncertainties.
@@
Figure 3.3: Comparison of our NNLO $F_{2}$ higher twist to that determined in protons and deuteron by Virchaux and Milsztajn in a NLO PDF fit to SLAC and BCDMS data [74]. They include TMCs, and use the same multiplicative parametrization of the HT, fitting proton and deuteron data separately, so the results are directly comparable to our own.
@@
When we add MHOUs to our NNLO fit, the HT correction is generally reduced in size. The same happens again when we go to aN3LO, where the MHOUs are significantly reduced. This shows that some of the supposed HT being seen at NNLO is in fact due to MHOUs, principally the N3LO perturbative correction to the leading twist contribution. This shows the difficulty in distinguishing phenomenologically between $Q^{2}$ dependence which is power like, and $Q^{2}$ dependence which is merely a power of a logarithm. This effect was also seen when going from NLO to NNLO in earlier attempts to determine HT [66]. One is forced to conclude that even the HT determined in the aN3LO fit may still be contaminated by a perturbative contribution. Similarly, $H_{2}^{d}$ may be tangled with the deuteron nuclear correction [31], and $H_{CC}$ with the nuclear corrections applied to much of the CC data [30].
@@ 3.2 Determination of Linear Power Corrections to Jets and Dijets
We now turn to the determination of power corrections for jet observables. The jet cross-sections included in the NNPDF4.0 fit are double differential single-inclusive jet data at $\sqrt{s}=8$ TeV and dijet cross-sections at $\sqrt{s}=7$ TeV from ATLAS [1, 2] and CMS [64, 53], following a detailed study of the reliability and impact of the datasets [8]. The jets are constructed in both cases using the anti-$k_{T}$ algorithm, and the data used in the fit were those with $R=0.6$ for ATLAS, $R=0.7$ for CMS.
The published datasets are at hadron level. Hadronization, underlying event and multi-parton corrections were estimated by the experimentalists from the spread in several runs of final state Monte Carlos with different tunes and input PDFs, was included in the covariance matrix. However the spread in the ATLAS estimates is very wide (consistent with zero), while that from CMS is much narrower, and significantly nonzero, particularly at low scale. Given this fact, it was decided in NNPDF4.0 not to correct the central values of the data, but to implement only an overall hadronization uncertainty as a contribution to the experimental covariance matrix. An internal cut on the jet $p_{T}$ was applied to the single-jet measurements of CMS in order to avoid points for which $p_{T}\lesssim 70$ GeV, where fixed-order calculations are not reliable. This cut is not necessary for ATLAS single-jet data since the lowest $p_{T}$ bin starts at $70$ GeV. No additional cuts are applied to dijet data, as the lowest value of the hard scale is one order of magnitude larger than that for the single-inclusive jet data. All these considerations will become relevant in the discussion below.
@@
In Fig. 3.4 we present the posterior distribution for the power corrections to single inclusive jets and dijets, and in Fig. 3.5 the related kinematic distribution of the data points. We observe two distinct patterns for single-inclusive jets and dijets. Starting with the single-inclusive jets, we find that not only is the prior significantly constrained, but also that the posterior distribution is peaked around $10$ GeV. Further, the posterior shows a weak dependence on the rapidity. It should be noted that the previous study in Ref. [70] estimated a positive power correction of about $7$ GeV for single-inclusive jet cross-sections, independent of rapidity. Our result is consistent with this estimate.
For the dijets, we find rather weaker evidence for a linear PC at central rapidity in the NNLO fit, with uncertainties remaining large. Note that for the dijet kinematics the bins with lowest invariant mass (and thus most likely to be affected by PCs) are at central rapidity, and even here the lowest invariant mass bin is still at a rather higher scale than the lowest $p_{T}$ scale of the single inclusive jet data.
@@
Figure 3.6: The hadronization correction as a function of $p_{T}$ for the 88 TeV single inclusive jet data, for $0<|y|<0.5$ (left) and $2.5<|y|<3.0$ (right). Our empirical estimates for the linear power correction at NNLO, NNLO+MHOU and N3LO+MHOU are compared to the ATLAS and CMS estimates obtained using final state Monte Carlo generators [1, 64].
@@
Finally, it is interesting to compare these empirical estimates of linear power corrections Eq. 2.25 with the estimates of hadronization, underlying event and multi-parton corrections made by ATLAS and CMS using final state Monte Carlos, which make no assumption regarding the $p_{T}$ dependence of the correction. The comparison for the CMS $8$ TeV single inclusive jet data [1, 64], for both central rapidity and high rapidity bins, is shown in Fig. 3.6. At central rapidity the agreement is very good, though at low $p_{T}$ the Monte Carlo estimates rise more steeply than our own, while at high $p_{T}$ they fall less quickly. At high rapidity the agreement is less good, the Monte Carlo estimate being rather higher, though still consistent within uncertainties. The consistency between these two very different approaches is quite encouraging, suggesting that if the Monte Carlo estimates of the hadronization (and multi-particle interactions) had been used to correct the data, converting it from hadron level to parton level, the linear power corrections found in our analysis would have been rather smaller, though more dependent on rapidity. Similar remarks apply to the dijet datasets.
@@
We evaluate the phenomenological consequences of including the HT corrections and PCs discussed in the previous sections at the level of a new global PDF fit, and discuss the impact on relevant LHC observables and on the determination of $\alpha_{s}$.
@@ 4.1 Global PDF fits with Power Corrections
The HT and power corrections discussed above were determined by performing global PDF fits to the NNPDF4.0 dataset, but lowering the kinematic cuts to increase sensitivity to the HT. However, we do not suggest performing a fit with these lowered cuts to determine the PDFs, since such a fit would include data which is significantly affected by HT. Instead, we use the HT and the PCs determined above to give corrections to the theory predictions in a fit with the conventional cuts used in NNPDF4.0 [34], i.e. $Q^{2}>3.49$ GeV2 and $W^{2}>12.5$ GeV2.
@@ 4.2 Fit Quality
We start by discussing the fit quality. In Tab. 4.1 we report the values of the experimental $\chi^{2}$ per data point for the global fit with and without HT and PCs, and for fits with perturbative calculations at NNLO, NNLO+MHOU, and aN3LO+MHOU. Together with the global $\chi^{2}$, values are also displayed for each separate process type. In order to disentangle the effect of the shifts in the theory predictions from the effect of the additional uncertainty due to the theory covariance matrix in eq. 4.2, we performed a series of fits where only the shifts are included in the theory predictions. The corresponding $\chi^{2}$ values are also reported in Tab. 4.1. All sources of correlations are fully accounted for, and the $t_{0}$ method [28] is used for multiplicative uncertainties. We refer to Ref. [37] for further details on the inclusion of the experimental uncertainties in the $\chi^{2}$.
@@ 4.3 PDFs and Parton Luminosities
In Fig. 4.3, the luminosities in the quark-quark, quark-antiquark, gluon-gluon, and gluon-quark channels are shown at $13$ TeV, as ratios of the fits with HT and PCs to the corresponding NNLO, NNLO+MHOU, and aN3LO+MHOU [37] standard fits. As expected from the PDF plots, the effects of HT and PCs at both high and low invariant masses are modest, well within one sigma, and largely disappear once MHOUs are added, and in the aN3LO fit. In fact the only really significant change due to the inclusion of PCs is in the gluon-gluon channel at invariant masses around the Higgs mass. This is particularly noticeable in the standard NNLO fit: here there is a one percent drop in the gluon-gluon luminosity for invariant masses between $100$ and $250$ GeV, just over two sigma. Once MHOUs are included the significance of this effect is reduced, and the aN3LO corrections reduce the impact to well within one sigma. A related effect is seen in the gluon-quark channel, though here it is never more than one sigma.
@@ 4.4 Predictions for LHC observables
We next analyse the impact of power corrections and higher twist on various totally inclusive LHC observables related to DY and Higgs production, specifically vector boson production, associated production of Higgs with a vector boson, then Higgs production in gluon fusion, and in vector boson fusion (VBF). The predictions use hard cross-sections computed up to N3LO using n3loxs [19] for associated production, ggHiggs [44] for gluon fusion, and proVBFH code [59] for VBF.
@@ 4.5 Effect on the Strong Coupling
We finally consider the effect of HT and PCs on the extraction of the strong coupling at the mass of the $Z$-boson, $\alpha_{s}(m_{Z})$, from the global dataset. We again use the theory covariance methodology, precisely as in Ref. [20], but now including also the HT and PC contributions to the theory covariance matrix. The results are shown in Fig. 4.6.
In the NNLO fit, without MHOU, the HT reduces $\alpha_{s}$ a little, while the jet PCs increase it, in each case by around one standard deviation. When MHOU is added, the best fit value of $\alpha_{s}$ drops significantly, and its uncertainty increases substantially, as found in Ref. [20]. Again the HT reduces $\alpha_{s}$, and the PCs increase it, by about one standard deviation. However, at aN3LO the pattern changes: while the aN3LO corrections reduce $\alpha_{s}$ a little, the HT correction now results in a substantial increase (more than a standard deviation). We think that this is because at aN3LO the HT is less contaminated by MHO perturbative corrections to the leading twist, and its effect on $\alpha_{s}$ is thus less easily hidden by the MHOU. The effect of the jet PCs is positive as before, but smaller (the jet cross-sections are still evaluated at NNLO+MHOU in the aN3LO fits). The result is a value of $\alpha_{s}(m_{Z})$ at aN3LO which is very consistent with that obtained at NNLO, but with substantially larger uncertainty.
As we noted in Ref. [20], all these results are subject to a further uncertainty due to the implementation of positivity constraints on the PDF fit. Since any PDF positivity constraint necessarily introduces a nongaussianity in the distribution of the PDFs, the closure testing of the methodology only works when the positivity constraints are removed. As can be seen in Fig. 4.6, releasing the positivity constraints results in lower values of $\alpha_{s}$ consistently across all the various determinations.